Peanuts worth 2.95/kg were mixed with cashews worth 6.25/kg to produce a mixture worth 5.00/kg. How many kilograms of each kind of nut were used to make 33kg of the mixture?
Answer by Paul(988)

2.95x+6.25(33-x)=5(33)
2.95x+206.25-6.25x=165
-3.3x=-41.25
x=12.5
33-12.5=20.5

Hence, for 2.95kg about 12.5KG of mixture is needed and for 6.25kg about 20.5KG of mixture is needed.
Paul.